[figure: 9.7 cm]

EQUATIONS OF MOTION FOR A CHARGED PARTICLE IN

n-DIMENSIONAL MAGNETIC FIELD

W. Florek and M. Thomas Institute of Physics, A. Mickiewicz Univ. ul. Umultowska 85, 61–614 Poznań, Poland

The equation of motion for a charged particle moving in the n-dimensional constant magnetic filed is obtained for any linear gauge and any metric tensor by generalization of Johnson and Lipmann’s approach. It allows to consider the magnetic orbits in the ndimensional space. It is shown that the movement of a particle can always be decomposed into a number of two-dimensional cyclotronic motions and a free particle part.

[figure: 13.4 cm]

Subject category :

6. Theory of Magnetism

Presentation mode : poster

Corresponding author : Wojciech Florek

Address for correspondence : Computational Phys. Div. Institute of Physics A. Mickiewicz Univ. ul. Umultowska 85 61–614 Poznań, Poland

Email address : florek@amu.edu.pl